# Measurement incompatibility in Bayesian multiparameter quantum estimation

### v1

@@ IV.1 Discrete quantum phase imaging
Consider a (d+1)(d{+}1)-mode interferometer, where $d$ hypotheses $\boldsymbol{\theta}$ for a set of unknown phases $\boldsymbol{\Theta}$ are encoded via the unitary operator $\mathrm{e}^{-iN^{j}\theta_{j}}$. Here, $N^{j}$ is the number operator of the $j$th mode, and the zeroth mode serves as a reference beam [jarzyna2012quantum]. This configuration models discrete quantum imaging and has been extensively explored in both local and Bayesian frameworks [macchiavello2003optimal, humphreys2013quantum, rubio2020bayesian, Albarelli2019c, gebhart2021bayesian]. In this context, a common probe state is the generalised N00N state
@@
To address this problem using the MSL, we restrict the hypothesis space $\Omega$ so that the phases are sufficiently localised to behave as location parameters, but remaining within a global regime. Specifically, we assume that, a priori, $\theta_{i}\in[-\pi/n,\pi/n]$, with $n\geq 4$, so that the largest prior width is $\pi/2$. For this configuration—and choosing a normalised $L$ with equal weights, i.e., $L=\mathbb{I}/d$—the average state, SPM operators, and SPM bound are [rubio2019quantum]
@@
which already exhibits Heisenberg scaling; this is consistent with the observations in Refs. [hall2012does, giovannetti2012subheisenberg, tsang2012zivzakai, pezze2013subheisenberg, rubio2018nonasymptotic].
@@
As such, the SPM bound is not tight, and we can thus expect incompatibility to arise in this system.
@@
However, it is straightforward to verify that $\mathcal{L}_{\mathrm{SPM}}\geq\mathcal{L}_{\mathrm{RPM}}$, consistent with the fact that $\mathcal{K}_{\mathrm{R}}$ in Eq. (69) is real-valued. Consequently, we shall instead evaluate the NH bound in Eq. (43) using the numerical algorithm of Ref. [multipar2025code] to obtain a tighter lower bound on the minimum MSL.
@@ IV.2 Phase and dephasing estimation
While phase estimation serves as the archetypal problem in quantum estimation, the simultaneous estimation of phase and phase diffusion plays an analogous role in multiparameter quantum estimation. The simplest single-qubit incarnation of this model exhibits measurement incompatibility in local QET [vidrighin2014joint], which however asymptotically disappears as many identical copies are measured collectively [roccia2018entangling]. This corresponds to preparing identical and factorised probe states, whereas asymptotically large entangled probes generally allow to optimally estimate both parameters [kobus202asymptotically]. Nevertheless, we consider here independent identical copies collectively measured, since this is a paradigmatic setting for investigating measurement incompatibility [vidrighin2014joint, roccia2018entangling, Yung2025].
For an initial state $\rho_{\mathrm{in}}$, the single-qubit phase and phase-diffusion channels [^6] The qubit can represent, e.g., a single-photon in optical interferometry, whereas the same setup for probe states with more photons is not equivalent to considering multiple copies of the single-photon state, but give rise to correlated dephasing [Genoni2011a, Szczykulska2017]., which commute, are defined as
@@
which mixes two uncertainty types [overton2024five]. Here, we assume $L=\mathbb{I}_{2}$. This corresponds to using the multiparameter symmetry function
@@
Following Ref. [overton2024five], Eq. (76) further leads to the ignorance prior [jaynes2003probability]
@@
Eq. (80) is the product of two independent priors—the first flat, representing maximum ignorance for a location, and the second Jeffreys’s prior, representing maximum ignorance for a scale [jaynes2003probability, linden2014bayesian]. The range for the dephasing parameter takes such a form when a symmetric, flat range of width $W_{2}$ is imposed in the log-transformed space.
Given this formulation, the state moments for a single copy can be calculated as
where $\operatorname{Ei}(z)$ denotes the exponential integral,
$\Xi(z)=\displaystyle\sideset{{}_{3}}{{}_{3}}{\mathop{F}}((1,1,1),(2,2,2),z)$ (83)
is a generalised hypergeometric function and $\gamma$ Euler’s constant. Although this is a qubit system amenable to further analytic treatment, the resulting expressions are not very informative. For our purposes, it suffices to note that the SPM operators arising from Eqs. (82) have the following structure:
@@
$S_{2}$ $\propto\mathbb{I}_{2}+h(W_{1},W_{2})\,\sigma_{x},$ (84b)
where $h$ is a function determined by the scalar coefficients appearing in the state operators defined above. As can be observed, the SPM operators do not generally commute, thereby guaranteeing the presence of measurement incompatibility at the single-copy level.
We next consider the $n$-copy state $\rho^{\otimes n}(\boldsymbol{\theta})$. Due to the prior, the corresponding $n$-copy average state and first moments, i.e., $\int\mathrm{d}\boldsymbol{\theta}\,p(\boldsymbol{\theta})\,\rho^{\otimes n}(\boldsymbol{\theta})$ and $\int\mathrm{d}\boldsymbol{\theta}\,p(\boldsymbol{\theta})\,\rho^{\otimes n}(\boldsymbol{\theta})\,\theta_{j}$, are not directly tensor powers of the single-copy $\rho_{0}$ and $\bar{\rho}_{j}$ in Eqs. (82).
In lieu of the PGM* bound (due to the high dimension of the $n$-copy state), we consider a local qubit tomography—where each copy is measured at random in the basis of one of the three Pauli operators. The outcomes are then processed with the optimal PM estimator, such that the resulting MSL lower bounds $\mathcal{L}_{\mathrm{prior}}$, which is the tighter bound than $\mathcal{L}_{\mathrm{PGM}}$ here.
Fig. 2 shows that the incompatibility window grows, with the incompatibility lower bound through the NH increasing, and the tomography upper bound increasing more rapidly. This goes against local QET results that, with collective measurements on multiple copies, anticipate the initially large incompatibility to decrease [vidrighin2014joint, roccia2018entangling]. However, this growth in incompatibility does come as $\mathcal{L}_{\mathrm{min}}$ and its upper and lower bounds themselves decrease in the number of copies. Furthermore, it is unclear that the $n$-copy $\mathcal{I}_{\mathrm{min}}$ necessarily drops, as $\mathcal{I}_{\mathrm{NH}}$ appears to level-off below the single-copy $\mathcal{I}_{\mathrm{tom}}$.
@@ IV.3 Qubit planar tomography
$\rho_{0}=\frac{\mathbb{I}}{2},\quad\bar{\rho}_{i}=\frac{v_{i}}{2}\sigma_{i}.$ (88)
Given that $\rho_{0}$ is proportional to the identity, the SPM operators are trivially obtained:
@@
The incompatibility bounds corresponding to the MSL upper and lower bounds discussed above are plotted in Fig. 3, for various choices of the prior parameters $W_{1}$, $W_{2}$ and $\alpha$. The graphical conventions distinguishing curves follow those established in Sec. IV.1. The region of incompatibility values permitted by the bounds is shaded in Fig. 3. We conjecture, however, that for this class of models the NH bound is attainable, i.e., $\mathcal{L}_{\mathrm{NH}}=\mathcal{L}_{\mathrm{min}}$, in analogy with local QET [Nagaoka1991].
@@ V Discussion
QET was initiated in the 1960s and 1970s, but only in recent years have several lines of research converged, leading to a deeper understanding of multiparameter aspects [demkowicz2020multi, Albarelli2019c, Pezze2025]. Building on this progress, in this work we have laid the foundations for a comprehensive theory of Bayesian multiparameter quantum estimation. We have achieved this by reviewing, streamlining, and contextualising existing results, and by deriving a number of new ones, including: the explicit set of conditions (19) that the optimal POVM must satisfy under the MSL criterion; the upper bound $2\mathcal{L}_{\mathrm{SPM}}$ on the minimum MSL in the first inequality of Eq. (35), showing that such a minimum can be at most twice the value of the bound when neglecting measurement incompatibility; the tighter upper bound based on the notion of PGM combined with the optimal PM estimator; the introduction of general PM operators and their corresponding matrix lower bounds via the monotonicity of inner products, presented in App. C; and finally, the demonstration that measurement incompatibility in a single-shot Bayesian setup is substantially affected by prior information, as illustrated through examples in qubit and optical configurations.
An important aspect that we have not addressed is the relationship between Bayesian QET and the bounds derived in local QET for the frequentist loss in Eq. (8). In this regard, we note that the authors of Ref. [demkowicz2020multi] conjectured that, under appropriate regularity conditions, the Bayesian MSL and the frequentist loss become asymptotically equivalent; that is,
@@
Since in local QET it holds that [Tsang2019, Carollo2019]
@@
where $\mathcal{L}_{\mathrm{S}}(\boldsymbol{\theta})$ is the Helstrom (SLD) CRB, our inequality in Eq. (35) is consistent with this asymptotic conjecture. Notice that, by definition, the Bayesian MSL $\mathcal{L}_{\mathrm{min}}^{(n)}$ is obtained by minimising over POVMs acting collectively on $n$ copies, dubbed “correlated measurements” in Ref. [Hayashi2023a].
Despite increasing efforts in the experimental implementation of few-copy collective measurements [roccia2018entangling, Hou2018, Yuan2020, Conlon2023, Mi2025, Yung2025], scaling them to large $n$ remains challenging in practice. Following this observation, recent work in local QET has addressed the minimum multiparameter frequentist loss achievable with single-copy measurements and classical feedback, referred to as “uncorrelated measurements” in Ref. [Hayashi2023a]. This quantity corresponds to a minimisation of the weighted trace of the inverse Fisher information matrix (FIM) over all single-copy POVMs,
@@
where $F(\boldsymbol{\theta})$ is the FIM obtained by applying a POVM to a single copy $\rho(\boldsymbol{\theta})$. This is often called the “most informative” [Nagaoka1989] or “tight” [Hayashi2023a] bound.
Unlike the HCRB, $\mathcal{L}_{\mathrm{min}}(\boldsymbol{\theta})$ can be larger than $2\mathcal{L}_{\mathrm{S}}(\boldsymbol{\theta})$ [Das2024, Hayashi2023a]. This appears to be in contrast with the Bayesian setting, where the inequality $2\mathcal{L}_{\mathrm{SPM}}\geq\mathcal{L}_{\mathrm{min}}$ holds at the single-copy level as well. The resolution of this apparent tension comes from noticing that the $\mathcal{L}_{\mathrm{min}}(\boldsymbol{\theta})$ is still attainable only asymptotically, but restricting each copy to be measured independently, albeit adaptively with classical feedback from previous outcomes. Therefore, we expect that imposing an analogous restriction in the Bayesian case should lead to a similar behaviour as the number of shots increases. More concretely, we conjecture that, under suitable regularity conditions
@@
Beyond the question of establishing formal connections with local QET, this work also opens several new avenues for future research, both theoretical and experimental; here we highlight a few.
@@
First, apart from measurement incompatibility, in multiparameter quantum metrology—i.e., in the simultaneous estimation of several parameters encoded by a quantum channel—additional intrinsic precision trade-offs may arise from the choice of probe state and control strategy [ragy2016compatibility, Gorecki2020, Hayashi2023a, Hu2024a], a phenomenon referred to as probe incompatibility [Albarelli2022]. While this aspect has not, to the best of our knowledge, been examined in detail within a Bayesian framework, the numerical approach for designing optimal protocols introduced in Ref. [Bavaresco2024designing] inherently accounts for probe incompatibility effects when extended to the multiparameter setting.
@@
In addition, while our analysis has focused on finite-dimensional systems, continuous-variable (CV) platforms—particularly quantum optical ones [demkowiczdobrzanski2015quantum, Polino2020, Pirandola2018, Barbieri2022]—play a central role in metrology and sensing applications. It will therefore be important to rigorously extend the theoretical framework developed here to CV systems. Given the rapid progress in learning theory for CV states [Wu2024g, Mele2024a, Bittel2025, Fanizza2025], this direction appears especially promising. In particular, the case of Gaussian states and priors [wang2007] may allow for a more tractable analysis while remaining highly relevant for practical implementations. In the single-parameter Bayesian setting, several specific results for Gaussian states are already known [Morelli2020], and a general framework for multiparameter local QET with Gaussian states has recently been proposed [Chang2025].
@@
Another important problem is understanding the effect of measurement—or other forms of—incompatibility for parameters obeying symmetries other than translation invariance over the real line, i.e., beyond location parameters. We have provided the groundwork to generalise the results of this study to any location-isomorphic estimation problem through the symmetry-informed mapping postulated in Eq. (76), concrete illustrations of which have been discussed both in Sec. IV.2 and in Ref. [overton2024five]. A more rigorous formulation should include the possibility of correlated priors and loss functions coupling different parameters, as well as the application of information geometry in a multidimensional parameter space [boeyens2021uninformed]. Moreover, a further extension to symmetries beyond location-isomorphic metrology is necessary to encompass phase parameters, which constitute a paradigmatic class requiring suitable trigonometric loss functions and priors [holevo2011probabilistic, Demkowicz-Dobrzanski2011].
@@
At the experimental level, our upper bound $2\mathcal{L}_{\mathrm{SPM}}$ suggests that the pseudo-gain $L^{ij}\operatorname{Tr}[\rho_{0}S_{j}S_{i}]$—i.e., the second term in the SPM lower bound (21); see App. A—offers a good compromise to use as the optimiser of experimental control parameters for individual shots in an adaptive protocol. This is due to its computational simplicity while also capturing much of the structure of the estimation problem at hand. Although such an approach will not generally lead to full optimality, we anticipate a substantial boost in quantum technological applications, in light of recent results from cold-atom experiments showing two-fold [glatthard2022optimal] and five-fold [overton2024five] precision enhancements obtained by employing the “classical” version of this quantifier, that is, the pseudo-gain computed using the optimal estimator and a fixed POVM. In this sense, $L^{ij}\operatorname{Tr}[\rho_{0}S_{j}S_{i}]$ may serve as a practical, global alternative to the weighted trace of the inverse quantum FIM.

### v2

@@ IV.1 Discrete quantum phase imaging
Consider a (d+1)(d{+}1)-mode interferometer, where $d$ hypotheses $\boldsymbol{\theta}$ for a set of unknown phases $\boldsymbol{\Theta}$ are encoded via the unitary operator $\mathrm{e}^{-iN^{j}\theta_{j}}$. Here, $N^{j}$ is the number operator of the $j$th mode, and the zeroth mode serves as a reference beam [140]. This configuration models discrete quantum imaging and has been extensively explored in both local and Bayesian frameworks [141, 38, 40, 65, 142]. In this context, a common probe state is the generalised N00N state
@@
To address this problem using the MSL, we restrict the hypothesis space $\Omega$ so that the phases are sufficiently localised to behave as location parameters, but remaining within a global regime. Specifically, we assume that, a priori, $\theta_{i}\in[-\pi/n,\pi/n]$, with $n\geq 4$, so that the largest prior width is $\pi/2$. For this configuration—and choosing a normalised $L$ with equal weights, i.e., $L=\mathbb{I}/d$—the average state, SPM operators, and SPM bound are [11]
@@
which already exhibits Heisenberg scaling; this is consistent with the observations in Refs. [143, 144, 10, 37, 145].
@@
As such, the SPM bound is not tight, and we can therefore expect incompatibility to arise in this system.
@@
However, it is straightforward to verify that $\mathcal{L}_{\mathrm{SPM}}\geq\mathcal{L}_{\mathrm{RPM}}$, consistent with the fact that $\mathcal{K}_{\mathrm{R}}$ in Eq. (69) is real-valued. Consequently, we shall instead evaluate the NH bound in Eq. (43) using the numerical algorithm of Ref. [111] to obtain a tighter lower bound on the minimum MSL.
@@ IV.2 Phase and dephasing estimation
While phase estimation serves as the archetypal problem in quantum estimation, the simultaneous estimation of phase and phase diffusion plays an analogous role in multiparameter quantum estimation. The simplest single-qubit incarnation of this model exhibits measurement incompatibility in local QET [146], which however asymptotically disappears as many identical copies are measured collectively [147]. This corresponds to preparing identical and factorised probe states, whereas asymptotically large entangled probes generally allow to optimally estimate both parameters [148]. Nevertheless, we consider here independent identical copies collectively measured, since this is a paradigmatic setting for investigating measurement incompatibility [146, 147, 149].
For an initial state $\rho_{\mathrm{in}}$, the single-qubit phase and phase-diffusion channels [^6] The qubit can represent, e.g., a single-photon in optical interferometry, whereas the same setup for probe states with more photons is not equivalent to considering multiple copies of the single-photon state, but give rise to correlated dephasing [150, 151]., which commute, are defined as
@@
which mixes two uncertainty types [35]. Here, we assume $L=\mathbb{I}_{2}$. This corresponds to using the multiparameter symmetry function
@@
Following Ref. [35], Eq. (76) further leads to the ignorance prior [44]
@@
Eq. (80) is the product of two independent priors—the first flat, representing maximum ignorance for a location, and the second Jeffreys’s prior, representing maximum ignorance for a scale [44, 113]. The range for the dephasing parameter takes such a form when a symmetric, flat range of width $W_{2}$ is imposed in the log-transformed space.
Given this formulation, the state moments for a single copy can be calculated to have form
$\rho_{0}$ $=\frac{1}{2}[\mathbb{I}_{2}+g(W_{1},W_{2})\,\sigma_{x}],$ (82a)
$\bar{\rho}_{1}$ $\propto\sigma_{y},$ (82b)
$\bar{\rho}_{2}$ $\propto\sigma_{x},$ (82c)
with the exact expressions given in App. D. Although this is a qubit system amenable to further analytic treatment, the resulting expressions are not very informative. For our purposes, it suffices to note that the SPM operators arising from Eqs. (145a) have the following structure:
@@
$S_{2}$ $\propto\mathbb{I}_{2}+h(W_{1},W_{2})\,\sigma_{x}.$ (83b)
As can be observed, the SPM operators do not generally commute, thereby guaranteeing the presence of measurement incompatibility at the single-copy level.
We next consider the $n$-copy state $\rho^{\otimes n}(\boldsymbol{\theta})$. Due to the prior, the corresponding $n$-copy average state and first moments, i.e., $\int\mathrm{d}\boldsymbol{\theta}\,p(\boldsymbol{\theta})\,\rho^{\otimes n}(\boldsymbol{\theta})$ and $\int\mathrm{d}\boldsymbol{\theta}\,p(\boldsymbol{\theta})\,\rho^{\otimes n}(\boldsymbol{\theta})\,\theta_{j}$, are not directly tensor powers of the single-copy $\rho_{0}$ and $\bar{\rho}_{j}$ in Eqs. (145a).
In lieu of the PGM bound (due to the high dimension of the $n$-copy state), we consider a local qubit tomography—where each copy is measured at random in the basis of one of the three Pauli operators. The outcomes are then processed with the optimal PM estimator, such that the resulting MSL lower bounds $\mathcal{L}_{\mathrm{prior}}$, which is again the tighter bound than $\mathcal{L}_{\mathrm{PGM}}$.
In Fig. 2 we present numerical results on measurement incompatibility, by evaluating upper and lower bounds as a function of the number of copies, for prior widths $W_{1}=\pi/2$ and $W_{2}=5$. The main plot shows that the incompatibility window grows with the number of copies, with the incompatibility NH lower bound increasing slightly, and the tomography upper bound increasing more rapidly. The latter result may be expected, since tomography corresponds to measuring each copy separately and thus cannot take advantage of collective effects. Interestingly, the overall behaviour appears to be qualitatively different from local QET, where the initially large incompatibility decreases by measuring multiple copies collectively [146, 147]. However, while incompatibility seemingly increases according to the upper and lower bounds, we emphasise that, given the large incompatibility window and that $\mathcal{I}_{\mathrm{NH}}$ appears to level-off below the single-copy $\mathcal{I}_{\mathrm{tom}}$, no definitive conclusion can yet be made regarding the behaviour of the actual $n$-copy $\mathcal{I}_{\mathrm{min}}$; a full characterisation of this behaviour remains an open task. Finally, we notice that this growth in incompatibility does come as the overall estimation precision increases, i.e., $\mathcal{L}_{\mathrm{min}}$ and its upper and lower bounds themselves decrease in the number of copies; see the plot of $\mathcal{L}_{\mathrm{SPM}}$ versus number of copies in the inset.
@@ IV.3 Qubit planar tomography
$\rho_{0}=\frac{\mathbb{I}}{2},\quad\bar{\rho}_{i}=\frac{v_{i}}{2}\sigma_{i},$ (87)
where for simplicity we label Pauli matrices as $\sigma_{1}\equiv\sigma_{x}$ and $\sigma_{2}\equiv\sigma_{y}$. Given that $\rho_{0}$ is proportional to the identity, the SPM operators are trivially obtained:
@@
The incompatibility bounds corresponding to the MSL upper and lower bounds discussed above are plotted in Fig. 3, for various choices of the prior parameters $W_{1}$, $W_{2}$ and $\alpha$. The graphical conventions distinguishing curves follow those established in Sec. IV.1. The region of incompatibility values permitted by the bounds is shaded in Fig. 3. We conjecture, however, that for this class of models the NH bound is attainable, i.e., $\mathcal{L}_{\mathrm{NH}}=\mathcal{L}_{\mathrm{min}}$, in analogy with local QET [86].
@@ V Discussion
One important aspect that we have not addressed is the relationship between Bayesian QET and the bounds derived in local QET for the frequentist loss in Eq. (8). In this regard, we recall a conjecture of Ref. [64]: under appropriate regularity conditions, the Bayesian MSL and the frequentist loss become asymptotically equivalent; that is,
@@
Since in local QET it holds that [76, 77]
@@
where $\mathcal{L}_{\mathrm{S}}(\boldsymbol{\theta})$ is the Helstrom (SLD) CRB, our inequality in Eq. (35) is consistent with this asymptotic conjecture. Notice that, by definition, the Bayesian MSL $\mathcal{L}_{\mathrm{min}}^{(n)}$ is obtained by minimising over POVMs acting collectively on $n$ copies, dubbed “correlated measurements” in Ref. [93].
Despite increasing efforts in the experimental implementation of few-copy collective measurements [147, 152, 153, 154, 155, 149], scaling them to large $n$ remains challenging in practice. Following this observation, recent work in local QET has addressed the minimum multiparameter frequentist loss achievable with single-copy measurements and classical feedback, referred to as “uncorrelated measurements” in Ref. [93]. This quantity corresponds to a minimisation of the weighted trace of the inverse Fisher information matrix (FIM) over all single-copy POVMs,
@@
where $F(\boldsymbol{\theta})$ is the FIM obtained by applying a POVM to a single copy $\rho(\boldsymbol{\theta})$. This is often called the “most informative” [85] or “tight” [93] bound.
Unlike the HCRB, $\mathcal{L}_{\mathrm{min}}(\boldsymbol{\theta})$ can be larger than $2\mathcal{L}_{\mathrm{S}}(\boldsymbol{\theta})$ [91, 93]. This appears to be in contrast with the Bayesian setting, where the inequality $2\mathcal{L}_{\mathrm{SPM}}\geq\mathcal{L}_{\mathrm{min}}$ holds at the single-copy level as well. The resolution of this apparent tension comes from noticing that the $\mathcal{L}_{\mathrm{min}}(\boldsymbol{\theta})$ is still attainable only asymptotically, but restricting each copy to be measured independently, albeit adaptively with classical feedback from previous outcomes. Therefore, we expect that imposing an analogous restriction in the Bayesian case should lead to a similar behaviour as the number of shots increases. More concretely, we conjecture that, under suitable regularity conditions
@@
We conclude by noting that hybrid figures of merit, incorporating methods from both local and global QET, have been introduced [19, 156]. However, these approaches are mostly heuristic, and understanding the connections outlined in this discussion could clarify their exact range of applicability.
VI Conclusions ad outlook
QET was initiated in the 1960s and 1970s, but only in recent years have several lines of research converged, leading to a deeper understanding of multiparameter aspects [64, 65, 67]. Building on this progress, in this work we have laid the foundations for a comprehensive theory of Bayesian multiparameter quantum estimation. We have achieved this by reviewing, streamlining, and contextualising existing results, and by deriving a number of new ones, including: the explicit set of conditions (19) that the optimal POVM must satisfy under the MSL criterion; the upper bound $2\mathcal{L}_{\mathrm{SPM}}$ on the minimum MSL in the first inequality of Eq. (35), showing that such a minimum can be at most twice the value of the bound when neglecting measurement incompatibility; the tighter upper bound based on the notion of PGM combined with the optimal PM estimator; the introduction of general PM operators and their corresponding matrix lower bounds via the monotonicity of inner products, presented in App. C; and finally, the demonstration that measurement incompatibility in a single-shot Bayesian setup is substantially affected by prior information, as illustrated through examples in qubit and optical configurations.
@@ VI Conclusions ad outlook
This work also opens several theoretical and experimental avenues for future research, a few of which we now highlight. For example, apart from measurement incompatibility, in multiparameter quantum metrology—i.e., in the simultaneous estimation of several parameters encoded by a quantum channel—additional intrinsic precision trade-offs may arise from the choice of probe state and control strategy [122, 157, 93, 124], a phenomenon referred to as probe incompatibility [123]. While this aspect has not, to the best of our knowledge, been examined in detail and in full generality within a Bayesian framework, numerical approaches for designing optimal multiparameter Bayesian estimation protocols [104, 105] inherently account for probe incompatibility effects.
@@
In addition, while our analysis has focused on finite-dimensional systems, continuous-variable (CV) platforms—particularly quantum optical ones [6, 2, 158, 159]—play a central role in metrology and sensing applications. It will therefore be important to rigorously extend the theoretical framework developed here to CV systems. Given the rapid progress in learning theory for CV states [160, 161, 162, 163], this direction appears especially promising. In particular, the case of Gaussian states and priors [101] may allow for a more tractable analysis while remaining highly relevant for practical implementations. In the single-parameter Bayesian setting, several specific results for Gaussian states are already known [164], and a general framework for multiparameter local QET with Gaussian states has recently been proposed [165].
@@
Another important problem is understanding the effect of incompatibility—measurement and other forms thereof—for parameters obeying symmetries other than translation invariance over the real line, i.e., beyond location parameters. We have provided the groundwork to generalise the results of this study to any location-isomorphic estimation problem through the symmetry-informed mapping postulated in Eq. (76), concrete illustrations of which have been discussed both in Sec. IV.2 and in Ref. [35]. A more rigorous formulation should include the possibility of correlated priors and loss functions coupling different parameters, as well as the application of information geometry in a multidimensional parameter space [117]. Moreover, a further extension to symmetries beyond location-isomorphic metrology is necessary to encompass phase parameters, which constitute a paradigmatic class requiring suitable trigonometric loss functions and priors [118, 166].
@@
At a practical level, identifying optimal measurements and controls for even an individual shot in an adaptive, experimental protocol can be computationally difficult. The upper bound $\mathcal{L}_{\mathrm{min}}\leq 2\mathcal{L}_{\mathrm{SPM}}$ suggests that optimising the SPM affords a good compromise between optimality and tractability, compared to the exact but difficult minimum loss, Eq. (20). This is due to its computational simplicity while also capturing much of the structure of the problem at hand, with a smaller SPM bound also meaning a smaller upperbound on $\mathcal{L}_{\mathrm{min}}$ and so the worst-case impact of measurement-incompatibility is smaller in absolute terms. In practice this sees the pseudo-gain $\operatorname{Tr}[\rho_{0}S_{j}S_{i}]$—i.e., the second term in the SPM lower bound (21); see App. A—become the optimiser of experimental control parameters for individual shots in an adaptive protocol, fulfilling the role of a global alternative to the weighted trace of the inverse quantum FIM. Although such an approach will not generally lead to full optimality, we anticipate a substantial boost in quantum technological applications, in light of recent results from cold-atom experiments showing two-fold [34] and five-fold [35] precision enhancements obtained by employing the “classical” version of this quantifier, that is, the pseudo-gain computed using the optimal estimator and a fixed POVM.
Altogether, we expect the Bayesian multiparameter framework for QET introduced here, together with the insight into the impact of measurement incompatibility, to become a fundamental building block for benchmarking both theory and experiment in quantum metrology, guiding the design of practically relevant, information–optimal quantum sensors.